Circular No. 7319

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gd IN NGC 2623 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), reports the discovery of an apparent supernova in NGC 2623. The supernova was discovered and confirmed on unfiltered images taken on Nov. 24.5 (mag ∼ 17.8) and 25.6 UT (mag ∼ 17.4), respectively. The new object is located at α = 8^{h}38^{m}24^{s}.61, δ = +25^{o}45′33′′.1 (equinox 2000.0), which is 7′′.3 east and 17′′.4 north of the nucleus of NGC 2623. KAIT images of the same field on Nov. 22.5 (limiting mag ∼ 18.0) and Nov. 5.5 (limiting mag ∼ 19.0) showed nothing at the position of the supernova.

SUPERNOVA 1999ga IN NGC 2442 E. P. Rubenstein, Cerro Tololo Interamerican Observatory, reports

V = 17.1 ± 0.1 from CCD images obtained at the 0.9-m telescope under photometric conditions on Nov. 25.1835 UT. He adds: “The B, R and I magnitudes were also measured to be 18.3±0.3, 16.2±0.1 and 15.6±0.1, respectively. It appears that this supernova is embedded in the galaxy’s dust lanes. In addition, the supernova’s red color suggests significant interstellar reddening. Spectroscopic follow-up should be made to determine if this is a type-II supernova, as is suggested by its location.”

HD 209458 Further to the note on IAUC 7317, G. Gonzalez points out that the

use of the two comparison stars permitted the calculation of an rms Δm for HD 209458 of 0.0065 mag prior to the start of the transit and that the magnitude difference between the comparison stars remained constant during the run to within the measurement errors. Ingress exterior contact was observed at the time stated, but the star’s low altitude did not allow detection of interior contact with confidence, and it is estimated that the dimming was no more than 0.028 mag. An 8-nm-wide filter was used.

NOTICE TO SUBSCRIBERS Owing to a change in the accounting procedures at the Smithsonian

Astrophysical Observatory, subscribers are advised that, beginning in December, payments should be made by mid-month, in order to guarantee their inclusion in the end-of-month accounting run.

1999 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden